OpenStudy (anonymous):

I dont get what to do For each function, (a) find f(0), (b) solve f(x). Heres the problems 1.f(x)=x^2-13x+40 2.f(x)=3-1/2x 3.f(x)=2-3x

OpenStudy (rulnick):

For the f(0), just replace x with 0. For instance: 1. f(0) = 0^2 -13(0) + 40 = 40.

OpenStudy (rulnick):

2. f(0)=3 3. f(0)=2

OpenStudy (rulnick):

For (b), there's missing info. Do you mean "solve f(x)=0"?

OpenStudy (anonymous):

yah sorry...

OpenStudy (rulnick):

For (2) is it 3-(1/2)x or 3-1/(2x)?

OpenStudy (anonymous):

IDK

OpenStudy (rulnick):

For (3) it is 0=2-3x, so x=2/3.

OpenStudy (rulnick):

For (1) it is 0 = x^2-13x+40, so x = 5 or x = 8 (two solutions).

OpenStudy (anonymous):

Thank you